OpenStudy (anonymous):

simplify the expression 144^/1/2

OpenStudy (driftracer305):

\[144^1/2\]??

OpenStudy (anonymous):

like 144 raised to -1/2

OpenStudy (driftracer305):

ohh negative... ok

OpenStudy (driftracer305):

so thats \[\frac{ 1 }{ \sqrt{144} }\]

OpenStudy (anonymous):

remember that \(144 = 12^2\)

OpenStudy (driftracer305):

@wio hey can i ask something?

OpenStudy (anonymous):

What is it?

OpenStudy (driftracer305):

do u knw who is a moderator here?

OpenStudy (anonymous):

Anyone with purple text name.

OpenStudy (driftracer305):

k... umm can they DELETE questions??

OpenStudy (anonymous):

For example, ganeshie8

OpenStudy (anonymous):

If it is against TOS they can.

OpenStudy (anonymous):

moderators are terrifying

OpenStudy (driftracer305):

ok @wio .......... thx a lot